OpenStudy (anonymous):

What is the solution to the equation (lny)^2 - 2ln(y^5) = 11, and how do you get the derivative of the equation with respect to y?

OpenStudy (anonymous):

wait, why are you trying to do a derivative of the equation? this is a simple quadratic equation.

OpenStudy (anonymous):

Like as in -b(+-) sqrt b^2-4ac over 2a?

OpenStudy (anonymous):

what is \[\ln a ^{b}\] ?

OpenStudy (anonymous):

Oh, so I can just plug all of that into the quadratic equation to get my solution and then take the derivative?

OpenStudy (anonymous):

No, you dont need to take any derivatives at all.

OpenStudy (anonymous):

I'm so confused.

OpenStudy (anonymous):

just answer my question. what is \[\ln a ^{b}\]

OpenStudy (anonymous):

a^b?

OpenStudy (anonymous):

okay do you know about logarithms?

OpenStudy (anonymous):

A little. I had calculus last semester...and I'm having to refresh everything...lol. If I took the derivative of the equation you just gave me, then I would get a^b right?

OpenStudy (anonymous):

I didn't ask for the derivative of log a^b. I just asked what log a^b was.

OpenStudy (anonymous):

okay log a^b = b log a

OpenStudy (anonymous):

Oh! Yes.

OpenStudy (anonymous):

do you know what logarithm means though?

OpenStudy (anonymous):

don't worry, thats a side question.

OpenStudy (anonymous):

Holy crap! So, really the problem should look like...\[((lny)^2)/(2lny^5)\]

OpenStudy (anonymous):

NO! how did you get THAT?

OpenStudy (anonymous):

(lny)^2 - 2lny^5 = 11 or (lny)^2 - 10 lny -11 = 0 say ln y = x x^2-10x-11 =0 solve for x, substitute in ln y

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

need feedback to know you got it.

OpenStudy (anonymous):

Ohhh OKay I got it, I'm sorry for the delay. I was making dinner. That makes sense. See, I started dividing because I figured if you have ln (2-3) then it's ln 2/ln 3. Idk. I don't know why I thought that.